# Global Dynamics of McKean-Vlasov SDEs via Stochastic Order

### v1

@@ 5.3 Local attractors of order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$
To investigate the local dynamics near the invariant measures of McKean-Vlasov SDEs, we establish a connecting orbit theorem and a structure theorem on local attractors of order-preserving semigroups.
Firstly, the following proposition concerns about connecting orbits on an order interval in $\mathcal{P}_{2}(\mathbb{R})$. It is a $\mathcal{P}_{2}$-version of the general connecting orbit theorem for an abstract monotone semiflow on a convex compact subset of a Hausdorff locally convex topological vector space (see Liu-Qu-Yao-Zhi [48, Theorem 4.2]).
Proposition 5.6.
Let $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ be an order-preserving semiflow, and suppose $\mu_{1}<_{\mathrm{st}}\mu_{2}$ are two order related invariant measures in $\mathcal{P}_{2}(\mathbb{R})$. Then at least one of the following holds:
  1. (a)
    $P_{t}^{*}$ has a third invariant measure $\nu\in\mathcal{P}_{2}(\mathbb{R})$ with $\mu_{1}<_{\mathrm{st}}\nu<_{\mathrm{st}}\mu_{2}$;
  2. (b)
    there exists an increasing connecting orbit $\{\nu_{t}\}_{t\in\mathbb{R}}$ from $\mu_{1}$ to $\mu_{2}$;
  3. (c)
    there exists a decreasing connecting orbit $\{\nu_{t}\}_{t\in\mathbb{R}}$ from $\mu_{2}$ to $\mu_{1}$.
Proof.
To prove the proposition, we only need to verify the assumptions (M1)-(M5) in [48, Theorem 4.2].
  1. (M1)
    Recall $(\mathcal{M}_{1}(\mathbb{R}),\left\lVert\cdot\right\rVert_{\mathcal{W}_{1}})$ is a normed space, see (5.10). A cone
    $C:=\bigg{\{}\mu\in\mathcal{M}_{1}(\mathbb{R}):\int_{\mathbb{R}^{d}}f\mathop{}\!\mathrm{d}\mu\geq 0\text{ for all nonnegative 1-Lipschitz increasing functions $f$}\bigg{\}}$
    in $\mathcal{M}_{1}(\mathbb{R})$ is given by [48, Lemma 3.5].
  2. (M2)
    By Proposition 3.4, we see that $S:=[\mu_{1},\mu_{2}]$ is compact in $\mathcal{P}_{2}(\mathbb{R})$, and hence, $S$ is compact in $\mathcal{P}_{1}(\mathbb{R})$. Recall $\mathcal{P}_{1}(\mathbb{R})$ is closed convex in $\mathcal{M}_{1}(\mathbb{R})$, see (5.11). Then $S$ is compact in $\mathcal{M}_{1}(\mathbb{R})$. The convexity and non-singleton of $S$ are obvious. By Corollary 3.5, (5.9) and the dual representation of $\mathcal{W}_{1}$, the 2-Wasserstein metric on $[\mu,\nu]$ induces the relative topology in $\mathcal{M}_{1}(\mathbb{R})$.
  3. (M3)
    By [48, Lemma 3.5], the partial order induced by the cone $C$ in the normed space $\mathcal{M}_{1}(\mathbb{R})$ coincides with the stochastic order when restricted on $\mathcal{P}_{2}(\mathbb{R})$. The infimum and the supremum of $S$ in $\mathcal{M}_{1}(\mathbb{R})$ are clearly endpoints $\mu_{1}$, $\mu_{2}$, of the order interval.
  4. (M4)
    The semiflow $P_{t}^{*}$ is order-preserving.
  5. (M5)
    $\mu_{1}$ and $\mu_{2}$ are invariant measures of $P^{\ast}_{t}$.
Based on the existence of connecting orbits, we will present a theorem for order-preserving semigroups that ensures the local attraction of the forward limit of a connecting orbit.
@@
it is obvious that $S$ is positively invariant. Since $B$ is positively invariant, we have $\omega(B)\subset\overline{B}$. Hence, $\mu_{0}$ is the unique invariant measure in $S$. Observing that $S$ has a lower bound $\nu^{-1,0}_{t_{-1}}\in S$ and an upper bound $\mu_{0}\in S$, Theorem 4.2 (i) (vi) and Remark 4.1 (i) entail that $\omega(S)=\{\mu_{0}\}$ is the global attractor of $P^{*}|_{S}$, where $P^{*}|_{S}$ means the semiflow $P^{*}$ restricted to $S$. The invariance of $\omega(B)$ and $\omega(B)\subset S$ entail that
@@
5.4 Proof of Theorem 2.3 and Corollary 2.4
We first prove Theorem 2.3 (i)-(iii). Then we use the structure theorems of global attractors (Theorem 4.2) and local attractors (Theorem 5.7) to show local convergence and attracting unbounded open sets (Theorem 2.3 (iv)(v)). To this end, we prepare two crucial propositions by tracing solutions backward along connecting orbits (see Proposition 5.8-5.9).
Proof of (i)-(iii) in Theorem 2.3.
(ii). By Theorem 5.4, we know that the open ball $B(\delta_{a_{k}},r_{k})$ is positively invariant under $P_{t}^{*}$, and there is some invariant measure $\nu_{2k-1}\in B(\delta_{a_{k}},r_{k})$. Next, we show that the open balls $\{B(\delta_{a_{k}},r_{k})\}_{k=1}^{n}$ are pairwise disjoint. Note that (2.1) implies
$r_{a_{i}}^{2}+r_{a_{j}}^{2}\leq 2^{-1}|a_{i}-a_{j}|^{2},\ \text{ for all }\ 1\leq i\neq j\leq n.$ (5.16)
Suppose on the contrary that there exists $\nu\in B(\delta_{a_{i}},r_{a_{i}})\cap B(\delta_{a_{j}},r_{a_{j}})$. Then,
$\mathcal{W}_{2}^{2}(\nu,\delta_{a_{i}})=\int_{\mathbb{R}}|x-a_{i}|^{2}\mathop{}\!\mathrm{d}\nu(x)\geq 2^{-1}|a_{i}-a_{j}|^{2}-\int_{\mathbb{R}}|x-a_{j}|^{2}\mathop{}\!\mathrm{d}\nu(x)=2^{-1}|a_{i}-a_{j}|^{2}-\mathcal{W}_{2}^{2}(\nu,\delta_{a_{j}}).$
Hence, $2^{-1}|a_{i}-a_{j}|^{2}\leq\mathcal{W}_{2}^{2}(\nu,\delta_{a_{i}})+\mathcal{W}_{2}^{2}(\nu,\delta_{a_{j}})<r_{a_{i}}^{2}+r_{a_{j}}^{2}$, which gives a contradiction to (5.16). This proves (ii) except (2.3), and (2.3) will be obtained later.
(i). For each $k=1,2,\dots,n-1$, notice $\frac{1}{2}(a_{k+1}-a_{k})^{2}\geq r_{k}^{2}+r_{k+1}^{2}\geq\frac{1}{2}(r_{k}+r_{k+1})^{2}$, so we have $r_{k}+r_{k+1}\leq a_{k+1}-a_{k}$. Together with $\nu_{2k-1}\in B(\delta_{a_{k}},r_{k})$ and $\nu_{2k+1}\in B(\delta_{a_{k+1}},r_{k+1})$, we have
$\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{2k-1}(x)<a_{k}+r_{k}\leq a_{k+1}-r_{k+1}<\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{2k+1}(x).$
Then Lemma 4.10 entails that $\nu_{1}<_{\mathrm{st}}\nu_{3}<_{\mathrm{st}}\cdots<_{\mathrm{st}}\nu_{2n-1}$. In the order interval $[\nu_{2k-1},\nu_{2k+1}]$, by Proposition 5.6, at least one of the three situations happens:
  • there is another invariant measure $\nu_{2k}$ with $\nu_{2k-1}<_{\mathrm{st}}\nu_{2k}<_{\mathrm{st}}\nu_{2k+1}$;
  • there is an increasing connecting orbit from $\nu_{2k-1}$ to $\nu_{2k+1}$;
  • there is a decreasing connecting orbit from $\nu_{2k+1}$ to $\nu_{2k-1}$.
We want to exclude the existence of connecting orbits, and we argue by contradiction. Assume there is an increasing connecting orbit from $\nu_{2k-1}$ to $\nu_{2k+1}$. Since $\nu_{2k-1}\in B(\delta_{a_{k}},r_{k})$ and $\nu_{2k+1}\in B(\delta_{a_{k+1}},r_{k+1})$, we have $P_{t_{0}}^{*}\mu\in B(\delta_{a_{k+1}},r_{k+1})$ for some $\mu\in B(\delta_{a_{k}},r_{k})$ and some $t_{0}>0$. But Theorem 5.4 and disjointness in (ii) tell us $P_{t}^{*}\mu\in B(\delta_{a_{k}},r_{k})$ for all $t\geq 0$, and $B(\delta_{a_{k}},r_{k})\cap B(\delta_{a_{k+1}},r_{k+1})=\emptyset$. This is a contradiction, so there is no increasing connecting orbit from $\nu_{2k-1}$ to $\nu_{2k+1}$. In the same manner, we can show there is no decreasing connecting orbit from $\nu_{2k+1}$ to $\nu_{2k-1}$. Hence, we obtain $(2n-1)$ order related invariant measures, say $\nu_{1}<_{\mathrm{st}}\nu_{2}<_{\mathrm{st}}\cdots<_{\mathrm{st}}\nu_{2n-1}$. Proposition 5.3 shows that there are no other invariant measures.
(ii)-(2.3) We only need to show that $\nu_{2k}\notin\overline{B(\delta_{a_{k}},r_{k})}$, and $\nu_{2k}\notin\overline{B(\delta_{a_{k+1}},r_{k+1})}$ can be proved in the same way. We first show that $\nu_{2k}\notin B(\delta_{a_{k}},r_{k})$. Assume otherwise, $\nu_{2k}\in B(\delta_{a_{k}},r_{k})$, and apply Proposition 5.6 to the order interval $[\nu_{2k},\nu_{2k+1}]$. Then at least one of the following statements holds:
  • there is another invariant measure $\mu$ with $\nu_{2k}<_{\mathrm{st}}\mu<_{\mathrm{st}}\nu_{2k+1}$;
  • there is an increasing connecting orbit from $\nu_{2k}$ to $\nu_{2k+1}$;
  • there is a decreasing connecting orbit from $\nu_{2k+1}$ to $\nu_{2k}$.
The statement (i) excludes the existence of $\mu$. Repeating the argument in the proof of (i), we can invoke Theorem 5.4 and disjointness in (ii) to show there are no connecting orbits, regardless increasing or decreasing. This is impossible, and thus, we must have $\nu_{2k}\notin B(\delta_{a_{k}},r_{k})$. By Remark 5.1, $\nu_{2k}\notin\partial\overline{B(\delta_{a_{k}},r_{k})}$, and hence, we have $\nu_{2k}\notin\overline{B(\delta_{a_{k}},r_{k})}$.
(iii). For each $k=1,2,\dots,n-1$, we prove there is a decreasing connecting orbit from $\nu_{2k}$ to $\nu_{2k-1}$, and the existence of an increasing connecting orbit from $\nu_{2k}$ to $\mu_{2k+1}$ can be shown similarly. In the order interval $[\nu_{2k-1},\nu_{2k}]$, we use again Proposition 5.6. Then at least one of the following three is true:
  • there is another invariant measure $\mu$ with $\nu_{2k-1}<_{\mathrm{st}}\mu<_{\mathrm{st}}\nu_{2k}$;
  • there is an increasing connecting orbit from $\nu_{2k-1}$ to $\nu_{2k}$;
  • there is a decreasing connecting orbit from $\nu_{2k}$ to $\nu_{2k-1}$.
The existence of $\mu$ is excluded by the statement (i). The statement (ii) suggests that $\nu_{2k-1}$ has an positively invariant neighbourhood $B(\delta_{a_{k}},r_{k})$. By (ii)-(2.3), $\nu_{2k}\notin\overline{B(\delta_{a_{k}},r_{k})}$. By repeating the proof in (i), an increasing connecting orbit does not exist. The only option left is a decreasing connecting orbit $\{\nu_{t}^{k,\downarrow}\}_{t\in\mathbb{R}}$ from $\nu_{2k}$ to $\nu_{2k-1}$. ∎
@@
Before the proof of (iv)(v) in Theorem 2.3, we need to verify a core condition in Theorem 5.7 that, $\omega(B(\delta_{a_{k}},r_{k}))$ has upper/lower bounds in connecting orbits. In Proposition 5.8-5.9, we will fix the connecting orbit as the laws of the solution flow of (1.1), which gives a time-inhomogeneous SDE. Noting that the solution of this SDE can be traced backward along the connecting orbit, and by the comparison theorem, we can find a measure in the connecting orbit at a point of time sufficiently close to $\infty$- ∞, which is a lower bound or upper bound of $\omega(B(\delta_{a_{k}},r_{k}))$ with respect to stochastic order.
@@
then there exists $t_{0}\in\mathbb{R}$ such that $\tilde{\nu}_{t}\leq_{\mathrm{st}}\omega(B)$ ((((resp. $\tilde{\nu}_{t}\geq_{\mathrm{st}}\omega(B)$)))) for all $t\leq t_{0}$.
@@
We first prove the “$<$” case in (5.17), and the case “$>$” can be proved similarly. Now consider the following time-inhomogeneous SDE
@@
Let $X_{t}^{s,\mu}$ be the unique solution of equation (5.18) at time $t$ with starting time $s$ and initial distribution $\mu$ and $P^{*}(t,s)$ be the dual semigroup of equation (5.18), i.e., $P^{*}(t,s)\mu:=\mathcal{L}(X_{t}^{s,\mu})$. Since $\{\tilde{\nu}_{t}\}_{t\in\mathbb{R}}$ is a connecting orbit, then Lemma 4.9 (i) yields that it is an entrance measure of equation (5.18), i.e., $P^{*}(t,s)\tilde{\nu}_{s}=\tilde{\nu}_{t}$ for all $t\geq s$. Note that $\sup_{t\in\mathbb{R}}\|\tilde{\nu}_{t}\|_{2}<\infty$ (by definition, the connecting orbit is bounded in $\mathcal{P}_{2}(\mathbb{R})$), by Lemma 4.9, we know that there exist $C>0,\lambda>0$ such that for any $\mu\in\mathcal{P}_{2}(\mathbb{R})$ and $t\geq s$,
@@
where $d_{2}$ is given as in (5.15). Note that
$\Big{|}\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\mathcal{L}(X_{t}^{s,\mu})(x)-\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\tilde{\nu}_{t}(x)\Big{|}\leq|f|_{\infty}d_{2}\big{(}\mathcal{L}(X_{t}^{s,\mu}),\tilde{\nu}_{t}\big{)},\ \text{ for all }\ f\in C_{b}(\mathbb{R}),$ (5.20)
and
$\Big{|}\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\mathcal{L}(X_{t}^{s,\mu})(x)-\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\tilde{\nu}_{t}(x)\Big{|}\leq d_{2}\big{(}\mathcal{L}(X_{t}^{s,\mu}),\tilde{\nu}_{t}\big{)}.$ (5.21)
By (5.19) and (5.20), we conclude that $\mathcal{L}(X_{t}^{s,\mu_{s}})\to\tilde{\nu}_{t}$ weakly as $s\to-\infty$ for all bounded sets $\{\mu_{s}\}_{s\leq t}$ in $\mathcal{P}_{2}(\mathbb{R})$. Then (5.19), (5.21) together with [68, Theorem 6.9] entail that, for all $t\in\mathbb{R}$,
@@
Next, we prove that $\tilde{\nu}_{t}\leq_{\mathrm{st}}\omega(B)$ for all $t\leq t_{0}$. By Theorem 2.1 (ii), the order interval $[\underline{\nu},\overline{\nu}]$ attracts the bounded set $B$. Then Proposition 3.4 and Lemma 4.1 (iii) tell us $\omega(B)$ is nonempty. For any $\rho\in\omega(B)$, it follows from Lemma 4.1 (i) that there exist $\{(t_{k},\rho_{k})\}_{k\geq 1}\subset\mathbb{R}^{+}\times B$ such that $t_{k}\to\infty$ as $k\to\infty$ and
@@
Note that $\rho_{k}\in B$ and $B$ is positively invariant, then Assumption 1 (iv), (5.23) and Theorem 4.3 yield
@@
Assume all assumptions in Theorem 2.3 hold. Let $\{\nu^{k,\uparrow}_{t}\}_{t\in\mathbb{R}}$ ((((resp. $\{\nu^{k,\downarrow}_{t}\}_{t\in\mathbb{R}}$)))) be the increasing ((((resp. decreasing)))) connecting orbit from $\nu_{2k}$ to $\nu_{2k+1}$ ((((resp. $\nu_{2k-1}$)))) in Theorem 2.3 (iii). Then there exists $t_{k}\in\mathbb{R}$ such that, for all $t\leq t_{k}$,
@@
there exists $\mu\in\overline{B(\delta_{a_{k+1}},r_{k+1})}$ such that $\int x\mathop{}\!\mathrm{d}\mu(x)=m_{2k}$. By the definition of $\Psi$ (see (5.13)) and Theorem 5.5, it follows that $\Psi(\mu)=\mu_{m_{2k}}=\nu_{2k}\in\Psi(\overline{B(\delta_{a_{k+1}},r_{k+1})})\subset B(\delta_{a_{k+1}},r_{k+1})$. This gives a contradiction to (2.3). Hence $m_{2k}\notin[a_{k+1}-r_{k+1},a_{k+1}+r_{k+1}]$.

### v2

@@ 4.2 Local attractors of order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$
To investigate the local dynamics near the invariant measures, we will present a theorem for order-preserving semigroups that ensures the local attraction of the forward limit of a connecting orbit. We first apply Proposition 4.3 and the main result of [36] to obtain the existence of connecting orbits.
Proof of (i)-(iii) in Theorem 1.4.
By Proposition 4.3, the equation (1.1) has at most (2n1)(2n-1) invariant measures. Then, by applying [36, Theorem 1.1], it remains only to show the pairwise disjointness of $\{B(\delta_{a_{k}},r_{k})\}_{k=1}^{n}$. In fact, (1.5) implies
$r_{i}+r_{j}\leq|a_{i}-a_{j}|,\ \text{ for all }\ 1\leq i\neq j\leq n.$ (4.3)
Suppose on the contrary that there exists $\nu\in B(\delta_{a_{i}},r_{i})\cap B(\delta_{a_{j}},r_{j})$. Then,
$|a_{i}-a_{j}|=\mathcal{W}_{2}(\delta_{a_{i}},\delta_{a_{j}})\leq\mathcal{W}_{2}(\delta_{a_{i}},\nu)+\mathcal{W}_{2}(\nu,\delta_{a_{j}})<r_{i}+r_{j},$
which gives a contradiction to (4.3). ∎
Based on the existence of connecting orbits, we give the following theorem ensuring the local attraction of the forward limit of a connecting orbit.
@@
it is obvious that $S$ is positively invariant. Since $B$ is positively invariant, we have $\omega(B)\subset\overline{B}$. Hence, $\mu_{0}$ is the unique invariant measure in $S$. Observing that $S$ has a lower bound $\nu^{-1,0}_{t_{-1}}\in S$ and an upper bound $\mu_{0}\in S$, Proposition 3.2 (i) (vi) and Remark 3.1 (i) entail that $\omega(S)=\{\mu_{0}\}$ is the global attractor of $P^{*}|_{S}$, where $P^{*}|_{S}$ means the semiflow $P^{*}$ restricted to $S$. The invariance of $\omega(B)$ and $\omega(B)\subset S$ entail that
@@
In the following, we will give key results to verify the core condition in Theorem 4.4 that, $\omega(B(\delta_{a_{k}},r_{k}))$ has upper/lower bounds in connecting orbits (Proposition 4.6 and Proposition 4.8). In Proposition 4.6 and Proposition 4.8, we will fix the connecting orbit as the laws of the solution flow of (1.1), which gives a time-inhomogeneous SDE. Noting that the solution of this SDE can be traced backward along the connecting orbit, and by the comparison theorem, we can find a measure in the connecting orbit at a point of time sufficiently close to -$\infty$, which is a lower bound or upper bound of $\omega(B(\delta_{a_{k}},r_{k}))$ with respect to stochastic order.
Firstly, we need to present a comparison principle for extrinsic-law-dependent SDEs (3.4).
Lemma 4.5.
Proof.
Since $\mathcal{L}(\xi)\leq_{\mathrm{st}}\mathcal{L}(\eta)$, Strassen’s theorem (see Lindvall [34, Theorem IV.2.4] or [35, equation (3)]) yields that there exist $\tilde{\xi},\tilde{\eta}\in L^{2}(\Omega,\mathcal{F}_{s},\mathbf{P})$ such that
$\mathcal{L}(\tilde{\xi})=\mathcal{L}(\xi),\ \ \mathcal{L}(\tilde{\eta})=\mathcal{L}(\eta),\ \ \text{ and }\ \ \tilde{\xi}\leq\tilde{\eta},\ \text{ $\mathbf{P}$-a.s.}$ (4.5)
[…]
Note that
Since $\theta>0$, according to the comparison theorem for usual SDEs [20, Theorem 1.2], we have
$X_{t}^{s,\mu,\tilde{\xi}}\leq Y^{s,\nu,\tilde{\eta}}_{t}\ \text{ for all }\ t\in[s,s+T],\text{ $\mathbf{P}$-a.s.}$
The desired result is proved. ∎
@@
then there exists $t_{0}\in\mathbb{R}$ such that $\tilde{\nu}_{t}\leq_{\mathrm{st}}\omega(B)$ ((resp. $\tilde{\nu}_{t}\geq_{\mathrm{st}}\omega(B)$)) for all $t\leq t_{0}$.
@@
We first prove the “<<” case in (4.7), and the case “>>” can be proved similarly. Now consider the following time-inhomogeneous SDE
@@
Let $X_{t}^{s,\mu}$ be the unique solution of equation (4.8) at time $t$ with starting time $s$ and initial distribution $\mu$ and $P^{*}(t,s)$ be the dual semigroup of equation (4.8), i.e., $P^{*}(t,s)\mu:=\mathcal{L}(X_{t}^{s,\mu})$. Since $\{\tilde{\nu}_{t}\}_{t\in\mathbb{R}}$ is a connecting orbit, then Lemma 3.7 (i) yields that it is an entrance measure of equation (4.8), i.e., $P^{*}(t,s)\tilde{\nu}_{s}=\tilde{\nu}_{t}$ for all $t\geq s$. Note that $\sup_{t\in\mathbb{R}}\|\tilde{\nu}_{t}\|_{2}<\infty$ (by definition, the connecting orbit is bounded in $\mathcal{P}_{2}(\mathbb{R})$), by Lemma 3.7, we have there exist $C>0,\lambda>0$ such that for any $\mu\in\mathcal{P}_{2}(\mathbb{R})$ and $t\geq s$,
@@
where $d_{2}$ is given by
$d_{2}(\nu_{1},\nu_{2})=\int_{\mathbb{R}}(1+|x|^{2})\mathop{}\!\mathrm{d}|\nu_{1}-\nu_{2}|(x).$ (4.10)
According to [54, Theorem 6.15], we conclude that
$\mathcal{W}_{2}\big(\mathcal{L}(X_{t}^{s,\mu}),\tilde{\nu}_{t}\big)\leq\sqrt{2d_{2}\big(\mathcal{L}(X_{t}^{s,\mu}),\tilde{\nu}_{t}\big)}\leq\sqrt{2C(1+\|\mu\|_{2}^{2})e^{-\lambda(t-s)}},\ \text{ for all }\ \mu\in\mathcal{P}_{2}(\mathbb{R}),\ t\geq s.$
Hence, for all bounded sets $\{\mu_{s}\}_{s\leq t}$ in $\mathcal{P}_{2}(\mathbb{R})$, we have
@@
Next, we prove that $\tilde{\nu}_{t}\leq_{\mathrm{st}}\omega(B)$ for all $t\leq t_{0}$. By Theorem 1.2 (ii), the order interval $\underline{\nu},\overline{\nu}]$ attracts the bounded set $B$. Then Proposition 2.4 and Lemma 3.1 (iii) tell us $\omega(B)$ is nonempty. For any $\rho\in\omega(B)$, it follows from Lemma 3.1 (i) that there exist $\{(t_{k},\rho_{k})\}_{k\geq 1}\subset\mathbb{R}^{+}\times B$ such that $t_{k}\to\infty$ as $k\to\infty$ and
@@
Note that $\rho_{k}\in B$ and $B$ is positively invariant, then Assumption 1 (iv), (4.12) and Lemma 4.5 yield
@@
Under Assumption 1, we have SDE (4.14) has a unique invariant measure $\rho^{\mu}\in\mathcal{P}_{2}(\mathbb{R})$ (see e.g., Meyn-Tweedie [41, 42] or Lemma 3.7 (ii)). Define a measure-iterating map $\Psi:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ by
[…]
According to [36, (67)], we have the following property for $\Psi$.
Lemma 4.7.
Remark 4.1.
Note that $\mu$ is an invariant measure of the equation (1.1) if and only if $\Psi(\mu)=\mu$. Then Lemma 4.7 implies that there is no invariant measure in $\partial\overline{B(\delta_{a},r_{a})}:=\overline{B(\delta_{a},r_{a})}\setminus B(\delta_{a},r_{a})$.
@@
Assume all assumptions in Theorem 1.4 hold. Let $\{\nu^{k,\uparrow}_{t}\}_{t\in\mathbb{R}}$ ((resp. $\{\nu^{k,\downarrow}_{t}\}_{t\in\mathbb{R}}$)) be the increasing ((resp. decreasing)) connecting orbit from $\nu_{2k}$ to $\nu_{2k+1}$ ((resp. $\nu_{2k-1}$)) in Theorem 1.4 (iii). Then there exists $t_{k}\in\mathbb{R}$ such that, for all $t\leq t_{k}$,
@@
there exists $\mu\in\overline{B(\delta_{a_{k+1}},r_{k+1})}$ such that $\int x\mathop{}\!\mathrm{d}\mu(x)=m_{2k}$. By the definition of $\Psi$ (see (4.15)) and Lemma 4.7, it follows that $\Psi(\mu)=\mu_{m_{2k}}=\nu_{2k}\in\Psi(\overline{B(\delta_{a_{k+1}},r_{k+1})})\subset B(\delta_{a_{k+1}},r_{k+1})$. This gives a contradiction to (1.7). Hence $m_{2k}\notin[a_{k+1}-r_{k+1},a_{k+1}+r_{k+1}]$.
@@
4.3 Proof of Main Theorems
We have obtained Theorem 1.2 and Corollary 1.3 in Section 3.3, and proved (i)–(iii) of Theorem 1.4 in Section 4.2. In this section, we complete the proof of the remaining main theorems and conclude with a conjecture on a saddle-point structure in the double-well granular media equation.